OpenStudy (anonymous):

Show that: L (\sigma\ i=1 to M c(n)*u(n)) = \sigma\ i=1 to M c(n)L(u(n))

OpenStudy (unklerhaukus):

\[L \left(\sum\limits_{i=1}^ M c(n)*u(n)\right) = \sum\limits_{ i=1}^M c(n)L\big(u(n)\big)\]? what does the L mean ?